OpenStudy (anonymous):

if log32=1.505 then what is the value of log_1000 32?

OpenStudy (anonymous):

\[\log _{1000}32 \] that the second equation

OpenStudy (anonymous):

im not sure what to do ??

OpenStudy (anonymous):

soooo did a little time ith my notes and i got .502 by the change of base rule

OpenStudy (anonymous):

is it correct??

OpenStudy (phi):

the first line means log base 10 (10 is understood) \[ \log 32 = \log_{10} 32 = 1.505 \] you could use the change of base formula: \[ \log_{new\ base} x = \frac{\log_{old\ base} x}{\log_{old\ base} \text{new base}}\]

OpenStudy (anonymous):

yay got it thank you thats what i used n___n @phi

OpenStudy (phi):

log(1000) = 3 so you should get \[ log_{1000} 32 = \frac{1.505}{3} = 0.5017 \approx 0.502\]

OpenStudy (anonymous):

yup thats it n_n thats what i got n_n thank you !! @phi

OpenStudy (phi):

if you use a calculator or type into google 1000^(0.5017) = you will get close to 32. (we would need more digits to get exactly 32)

OpenStudy (anonymous):

ohhhh ok then